The fixed cost of a company is 35000 and the variable cost per unit is 500 and the revenue function for sale is X units is given by R(x)=5000X-100X2

I) find the profit function

Ii) break even values

III)the values of X which results in a loss


1
Expert's answer
2022-06-16T15:02:41-0400

(i)

Given "\text{TFC of product} = 35\,000", "\text{TVC} = 500x", where "x=\text{no. of units of product}", "\text{required cost function} = C(x)=35\,000+500x".

Given "R(x)=5\,000x-100x^2".

Thus "P(x)=R(x)-C(x)=5\,000x-100x^2-35\,000-500x",

"P(x)=4\,500x-100x^2-35\,000"

(ii)

For breakeven point, "P(x)=0" "\Rightarrow" "4\,500x-100x^2-35\,000=0" "\Rightarrow" "x=10, x=35".


(iii)

For loss, "P(x)<0" "\Rightarrow" "x^2-45x+350>0" "\Rightarrow" "x>35" or "x<10".